OpenStudy (anonymous):

determine whether y varies directly with x, if so find the constant of variation k. 8y=7x-27

OpenStudy (anonymous):

hhhhhhhhhhheeeeeeeeeeeeelllllllllllllllllllllpppppppppppp me please

OpenStudy (anonymous):

haha always me.

OpenStudy (anonymous):

yea thanks

OpenStudy (anonymous):

ok i believe it does but i dont understand the second part of the question.

OpenStudy (anonymous):

i really dont get it and this is algebra 2 so the options it gives me are as follows: yes;8, yes;7, yes;7/8,or no

OpenStudy (lgbasallote):

@hereweg0 how did you know it does with that kind of info :o you must be veeerrryyyy good

OpenStudy (anonymous):

i remember some parts but i dont remember all... thats the prob with me right now.

OpenStudy (anonymous):

and i have another question with that and it is a chart

OpenStudy (anonymous):

8y=7x-27 solving for y y = 7x / 8 -27 / 8 y = 7/8 (x - 27/ 7) considering 7/8 as constant K y = K ( x - 27 / 7) which indicates y ∞ (x - 27 / 7) indicating that y varies directly with x and value of K is 7 / 8

OpenStudy (anonymous):

i dont kno is that wat i have to do

OpenStudy (anonymous):

imnot so sure either.

OpenStudy (anonymous):

np